A Complete Bibliography of ACM/IMS Transactions on Data Science (TDS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 August 2023 Version 1.03

Title word cross-reference

3 [ZLJC21]. K [AAUQ21].

-Subset [AAUQ21].

Abstractive [SKRR21]. across [GLZY20]. Active [NGMQ20]. Activity [YZL^{+}20]. Adaptive [XFW^{+}20]. Addition [YLY20]. Adversarial [DLZ^{+}21]. Air [LNZ21]. Algorithm [WLM20, ZCA21]. Algorithmic [CDK^{+}21]. Allocation [DZWL20]. Analysis [CPAS20, CAG20, LHJ20, TLL^{+}20]. Analytics [DZWL20, ILI^{+}21, LNP21]. Anomaly [AIR21, ZCCB21]. Anonymizing [WLM20]. Apache [ZLZ^{+}20]. Application [JWK^{+}21]. Applications [AAUQ21, BL20, DZWL20, ZS21]. Approach [ILI^{+}21, SJ20]. Approximate [ZMS^{+}21]. Assessment [LWG^{+}21]. Association [MPS^{+}21]. Asterisk [NGMQ20]. Attention [CGS^{+}21]. Attributes [TZW^{+}21]. Augmentation [GSM21]. Automatic [NGMQ20]. Autonomous [BCB^{+}21]. aware [LZYZ21, LWG^{+}21, XGB^{+}21].

Backup [YDZH21]. Bandits [AAUQ21]. Based [FZL^{+}20, LCZ^{+}20, LZYZ21, XGB^{+}21, ZLJC21, BVS^{+}20, GSM21, LWG^{+}21, WGL^{+}21, ZS21]. Behavior [TLL^{+}20]. Belief [LNZ21]. Bias [XGB^{+}21]. Bias-aware [XGB^{+}21]. Big [DZWL20, SHC^{+}21, TLL^{+}20, XGB^{+}21, ZMS^{+}21].

Blackmarket [DC21b]. Blackmarket-Driven [DC21b]. board [LKW^{+}20]. Boosted [YLY20]. Boosting [YDZH21]. Buildings [ILI^{+}21]. Bus [AL20].

Card [TZW^{+}21]. CBIR [MB21]. Centrality [DBR21]. Characteristics [FZL^{+}20]. Cities [JSF^{+}21, LBZB20, LBZB21, XS20]. Classification [ZLJC21]. Classifying [YDZH21]. Clinical [DC21a]. CNNs [ZLJC21]. Collaborative [SS20]. Collusion [DC21b]. Communities [CPAS20]. complete [SHC^{+}21]. Complex [BCC^{+}21, CPAS20]. Component [CAG20]. Computing [LBZB20, LBZB21]. Context [AL20]. Context-specific [AL20]. Continuous [BCC^{+}21]. coSense [SS20]. Coupled [BCC^{+}21]. Crowding [AL20]. Current [ZCCB21]. Cyberbullying [CGS^{+}21].

D [ZLJC21]. D-CNNs [ZLJC21]. Dark [LWG^{+}21]. Data [AIR21, AL20, DC21a, DZWL20, GSM21, HSC20, ILI^{+}21, LNP21, LKW^{+}20, LLC^{+}20, LWG^{+}21, NGMQ21, SHC^{+}21, SGJ^{+}21, TLL^{+}20, TZW^{+}21, WLM20, XGB^{+}21, YDZH21, ZMS^{+}21, ZS21]. Data-driven [AL20, ILI^{+}21]. Datasets [NGMQ20]. DataStorm [BCC^{+}21]. Deduplication [YDZH21]. Deep [DCZ^{+}21, GLD^{+}21, LHJ20, LWG^{+}21, MB21, SJ20, WGL^{+}21]. Derived [MB21]. Descent [DCZ^{+}21]. Detect [ZCCB21]. Detection [CPAS20, CGS^{+}21, NGMQ21, XS20]. Detectors [ZCCB21]. Developments [DC21a]. Diagnostics [LKW^{+}20]. Differentially [DCZ^{+}21]. Disaggregation [DLZ^{+}21]. Discovery [BFDS20, JWK^{+}21, WC20]. Distance [GSM21]. Distance-based [GSM21]. Driven [DC21b, AL20, ILI^{+}21]. Dynamic [FZL^{+}20].

Edge [YLY20]. Editorial [Ooi20]. Educational [XGB^{+}21]. Effective [BFDS20]. Efficiency [ILI^{+}21]. Efficient [XS20, ZLZ^{+}20, LKW^{+}20]. Embedding [JSF^{+}21]. Embeddings [TSS20]. Employment [XGB^{+}21]. Empty [BCB^{+}21]. Energy [DLZ^{+}21, ILI^{+}21, LKW^{+}20]. Energy-efficient [LKW^{+}20]. Ensemble [CPAS20]. Entity [TLL^{+}20]. Environments [BCC^{+}21]. Erroneous [NGMQ21]. Estate [XFW^{+}20]. Event [BVS^{+}20, LNP21]. Exemplar [YZL^{+}20]. Explaining [TSS20].

Fairness [EAYG20]. Feature

[XFW^{+}20, ZLJC21]. Feature-Influence [XFW^{+}20]. Features [AL20, MB21, SJ20]. Feedback [AAUQ21]. Female [YXL20]. Films [YXL20]. Financial [TSS20]. Fine [LHJ20]. Fine-Tuning [LHJ20]. Finite [XFW^{+}20]. Framework [WGL^{+}21, XS20].

Generating [NGMQ20]. Genetic [ZCA21]. GeoMatch [ZLZ^{+}20]. GLOVE [GFFS21]. Gradient [DCZ^{+}21]. Graduate [XGB^{+}21]. granularity [GLD^{+}21]. Graph [WGL^{+}21, ZMS^{+}21]. Graphs [LZYZ21]. Group [LCZ^{+}20]. Group-Based [LCZ^{+}20]. Guided [JWK^{+}21, LKW^{+}20].

HandiText [FZL^{+}20]. Handwriting [FZL^{+}20]. Hash [LWG^{+}21]. Hash-based [LWG^{+}21]. Hierarchical [CGS^{+}21, ZMS^{+}21]. Hotspot [XS20]. Human [JSF^{+}21]. Hybrid [DZWL20].

Identification [BL20]. Identifying [XFW^{+}20]. Image [LWG^{+}21, ZLJC21]. Imagery [SJ20]. Impact [YXL20]. Inaugural [Ooi20]. Incremental [FZL^{+}20]. Individualized [GLZY20]. Indoor [LLC^{+}20]. Influence [AAUQ21, XFW^{+}20, YLY20]. Information [DBR21]. Interactive [WGL^{+}21]. Internet [HSC20, SS20]. Internet-of-Things [SS20]. Interval [BVS^{+}20]. Interval-based [BVS^{+}20]. Introduction [HSC20, LBZB20, LBZB21, ZS21]. Intrusions [ZCCB21]. IoT [BVS^{+}20, DZWL20]. Irregular [GLD^{+}21]. Issue [HSC20, LBZB20, LBZB21, Ooi20, ZS21]. Iterative [DCZ^{+}21].

Job [EAYG20].

kD [SGJ^{+}21]. kD-STR [SGJ^{+}21]. Knowledge [LZYZ21].

Lab [ZCA21]. Lake [JWK^{+}21]. Landscape [ZCCB21]. Large [AIR21, DBR21, NGMQ20, WLM20, ZLZ^{+}20]. Large-Scale [AIR21, WLM20, ZLZ^{+}20]. Learning [DCZ^{+}21, HSC20, JWK^{+}21, MB21, NGMQ21, SJ20, WGL^{+}21, ZLJC21, ZS21]. Learning-based [ZS21]. Level [GFFS21]. Linear [AAUQ21]. Linguistic [LZYZ21]. Local [GLD^{+}21]. Lossy [BVS^{+}20]. LSTM [FZL^{+}20].

Machine [JWK^{+}21]. Map [ZLZ^{+}20]. Mapping [SJ20]. Marketplaces [EAYG20]. Matching [ZLZ^{+}20]. Maximization [AAUQ21, GCY^{+}21]. Maximizing [YLY20]. Meaningful [BFDS20]. Measuring [YXL20]. Media [BL20, DC21b]. Metadata [YDZH21]. Method [SGJ^{+}21]. Metro [TZW^{+}21]. Middleware [SS20]. Mining [DC21a, GLZY20]. Minority [GSM21]. Mitigation [BL20]. Mixture [XFW^{+}20]. Mobile [GFFS21, LNP21]. Mobility [JSF^{+}21, LLC^{+}20]. Model [LHJ20, XFW^{+}20]. Modeling [CGS^{+}21, LNZ21]. Modelling [SGJ^{+}21]. Models [AL20, SKRR21]. Multi [GLZY20, GLD^{+}21, ZCA21]. Multi-granularity [GLD^{+}21].

Multi-Sequences [GLZY20]. Multi-user [ZCA21]. Multifaceted [LHJ20]. Multiple [JSF^{+}21]. Multivariate [WC20].

Nearest [ZMS^{+}21]. Necessary [CDK^{+}21]. Neighbor [ZMS^{+}21]. Network [DBR21, GYZ^{+}21]. Networks [CPAS20, CGS^{+}21, LCZ^{+}20, YLY20]. Neural [GYZ^{+}21, LCZ^{+}20, LZYZ21, SKRR21]. Nodes [DBR21]. Non [AAUQ21]. Non-Linear [AAUQ21]. Nondominated [ZCA21].

On-board [LKW^{+}20]. Online

[DC21b, EAYG20, YLY20]. Optimization [DCZ^{+}21]. Outlier [BFDS20]. Oversampling [GSM21].

Parallel [WLM20]. Path [LKW^{+}20].

Patterns [CGS^{+}21, WC20]. Performance [YDZH21]. Phone [GFFS21, LNP21]. Physics [JWK^{+}21, LKW^{+}20]. Physics-Guided [JWK^{+}21, LKW^{+}20]. Placement [SHC^{+}21]. PoBery [SHC^{+}21]. POI [JSF^{+}21, LCZ^{+}20]. Popular [GCY^{+}21]. Positioning [LLC^{+}20]. Possible [CDK^{+}21]. Possibly [SHC^{+}21]. Possibly-complete [SHC^{+}21]. Prediction [AL20, LNZ21, XGB^{+}21]. Preferences [TZW^{+}21]. Preserving [AIR21, GFFS21, DC21a]. Principal [CAG20]. Privacy [BL20, DC21a, GFFS21]. Privacy-Preserving [GFFS21, DC21a]. Private [DCZ^{+}21]. Probabilistic [GSM21, SHC^{+}21]. Products [GCY^{+}21]. Profiles [JWK^{+}21]. Profitable [GCY^{+}21]. Propagation [DBR21]. Publishing [GFFS21].

Quality [LWG^{+}21, LNZ21]. Quantized [GYZ^{+}21]. Queries [BVS^{+}20, SHC^{+}21].

Raw [LLC^{+}20]. Real [CAG20, XFW^{+}20].

Real-Time [CAG20]. Recognition [FZL^{+}20, GLD^{+}21, LHJ20]. Recommendation [LCZ^{+}20, WGL^{+}21]. Reconstruction [ZLJC21]. Record [GFFS21]. Record-Level-Truthful [GFFS21]. Recurrent [LCZ^{+}20]. Redistribution [BCB^{+}21]. Reduction [SGJ^{+}21]. Reinforcement [WGL^{+}21]. Relationships [BFDS20]. Relevance [LWG^{+}21]. Relevance-aware [LWG^{+}21]. Remote [ZCA21]. Representation [NGMQ21, WGL^{+}21, YXL20]. Representations [GLD^{+}21]. Restoring [YDZH21]. Retrieving [HSC20]. Revenue [GCY^{+}21]. Road [SJ20]. Robust [XS20]. Role [DBR21]. Route [TZW^{+}21]. Rules [MPS^{+}21].

Safety [SJ20]. Sampling [AIR21]. Satellite [ZMS^{+}21]. Scalable [LNZ21]. Scale [AIR21, DBR21, ILI^{+}21, WLM20, ZLZ^{+}20]. Scanning [SHC^{+}21]. Scene [GLD^{+}21]. Scheduling [ZCA21]. Science [ZS21]. Scientific [JWK^{+}21]. Scoring [EAYG20]. Search [ZMS^{+}21]. Searching [YZL^{+}20]. Seed [DBR21]. Selection [LKW^{+}20]. Self [WGL^{+}21]. Self-supervised [WGL^{+}21]. Semantics [AIR21, LLC^{+}20]. Sensing [SS20]. Sequence [SKRR21]. Sequence-to-Sequence [SKRR21]. Sequences [GLZY20]. Series [AIR21, WC20]. Set [ZLJC21]. Set-Based [ZLJC21]. Simplex [ZCA21]. Simulating [JWK^{+}21]. Simulations [BCC^{+}21]. Simultaneous [ZLJC21]. Smart [LBZB20, LBZB21, TZW^{+}21, XS20]. Social [AAUQ21, BL20, YLY20]. Sorting [ZCA21]. Space [AAUQ21]. Spark [ZLZ^{+}20]. Spatial [WC20]. Spatio [SGJ^{+}21, WC20]. Spatio-Temporal [SGJ^{+}21, WC20]. Special [HSC20, LBZB20, LBZB21, ZS21]. Specialized [TSS20]. specific [AL20]. Spread [YLY20]. STAR [MPS^{+}21]. Steganography [LZYZ21]. Stochastic [AAUQ21]. STR [SGJ^{+}21]. Strategy [AIR21]. Streams [BVS^{+}20]. Streetview [SJ20]. Study [TZW^{+}21]. Subset [AAUQ21]. Summarization [SKRR21]. Summarizing [MPS^{+}21]. SUMMER [XGB^{+}21]. supervised [WGL^{+}21]. Supervision [NGMQ20]. Support [ZS21]. Survey [BL20, DBR21, DC21b]. Synthetic [GSM21]. System [ZMS^{+}21]. Systems [BCB^{+}21, TZW^{+}21].

TabReformer [NGMQ21]. Task [DZWL20]. Taxi [BCB^{+}21]. Techniques [CDK^{+}21]. Temperature [JWK^{+}21]. Temporal [CGS^{+}21, SGJ^{+}21, WC20]. Tensor [GYZ^{+}21]. Text [GLD^{+}21, SKRR21, WGL^{+}21]. Text-based [WGL^{+}21]. Things [SS20, HSC20]. Threat [ZCCB21]. Thresholds [GLZY20]. Time [AIR21, BCB^{+}21, CAG20, TZW^{+}21, WC20, YXL20]. Timed [MPS^{+}21]. Timetable [ZCA21]. Top [AAUQ21]. Topic [LZYZ21]. Topic-aware [LZYZ21]. Tourism [LNP21]. Training [NGMQ20]. Trajectories [GFFS21, YZL^{+}20]. Trajectory [WLM20]. Transfer [JSF^{+}21]. Translating [LLC^{+}20]. Travel [TZW^{+}21]. TRIPDECODER [TZW^{+}21]. Truthful [GFFS21]. Tuning [LHJ20].

Uncertainty [TSS20]. Unified [XS20]. Unsupervised [NGMQ21, ZCCB21]. Updating [LNZ21]. Urban [BCC^{+}21, DZWL20, JSF^{+}21, LBZB20, LBZB21, LNZ21, TLL^{+}20]. User [TLL^{+}20, ZCA21]. Using [AL20, LKW^{+}20, MB21, WGL^{+}21, ZCA21]. Utility [GLZY20].

Vehicle [BCB^{+}21]. via [JSF^{+}21]. Visual [LHJ20].

WattScale [ILI^{+}21]. Weighted [XFW^{+}20]. Windows [BCB^{+}21]. Winners [CDK^{+}21].

Word [TSS20].

References

Agarwal:2021:STS

[AAUQ21] Mridul Agarwal, Vaneet Aggarwal, Abhishek K. Umrawal, and Christopher J. Quinn. Stochastic top K-subset bandits with linear space and non-linear feedback with applications to social influence maximization. ACM Transactions on Data Science (TDS), 2(4):38:1–38:39, November 2021. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3507787.

Ahmed:2021:SAP

[AIR21] Shibbir Ahmed, Md Johirul Islam, and Hridesh Rajan. Semantics and anomaly preserving sampling strategy for large-scale time series data. ACM Transactions on Data Science (TDS), 2(4):41:1–41:??, November 2021. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3511918.

Arabghalizi:2020:DDB

[AL20] Tahereh Arabghalizi and Alexandros Labrinidis. Data-driven bus crowding prediction models using context-specific features. ACM Transactions on Data Science (TDS), 1(3):23:1–23:33, October 2020. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3406962.

Babicheva:2021:EVR

[BCB^{+}21] Tatiana Babicheva, Matej Cebecauer, Dominique Barth, Wilco Burghout, and Leïla Kloul. Empty vehicle redistribution with time windows in autonomous taxi systems. ACM Transactions on Data Science (TDS), 2 (1):5:1–5:22, January 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3416915.

Behrens:2021:DCC

[BCC^{+}21] Hans Walter Behrens, K. Selçuk Candan, Xilun Chen, Yash Garg, Mao-Lin Li, Xinsheng Li, Sicong Liu, Maria Luisa Sapino, Md Shadab, Dalton Turner, and Magesh Vijayakumaren. DataStorm: Coupled, continuous simulations for complex urban environments. ACM Transactions on Data Science (TDS), 2(3): 19:1–19:37, August 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3447572.

Bessa:2020:EDM

[BFDS20] Aline Bessa, Juliana Freire, Tamraparni Dasu, and Divesh Srivastava. Effective discovery of meaningful outlier relationships. ACM Transactions on Data Science (TDS), 1(2):12:1–12:33, July 2020. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3385192.

Beigi:2020:SPS

[BL20] Ghazaleh Beigi and Huan Liu. A survey on privacy in social media: Identification, mitigation, and applications. ACM Transactions on Data Science (TDS), 1 (1):7:1–7:38, March 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
abs/10.1145/3343038.

Busany:2020:IBQ

[BVS^{+}20] Nimrod Busany, Han Van Der Aa, Arik Senderovich, Avigdor Gal, and Matthias Weidlich. Intervalbased queries over lossy IoT event streams. ACM Transactions on Data Science (TDS), 1(4):27:1– 27:27, December 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3385191.

Chowdhury:2020:RTP

[CAG20] Ranak Roy Chowdhury, Muhammad Abdullah Adnan, and Rajesh K. Gupta. Real-time principal component analysis. ACM Transactions on Data Science (TDS), 1(2):10:1–10:36, July 2020. CODEN ???? ISSN 2691-

1922. URL https://dl.acm.
org/doi/10.1145/3374750.

Chakraborty:2021:ATN

[CDK^{+}21] Vishal Chakraborty, Theo Delemazure, Benny Kimelfeld, Phokion G. Kolaitis, Kunal Relia, and Julia Stoyanovich. Algorithmic techniques for necessary and possible winners. ACM Transactions on Data Science (TDS), 2(3): 22:1–22:23, August 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3458472.

Cheng:2021:MTP

[CGS^{+}21] Lu Cheng, Ruocheng Guo, Yasin N. Silva, Deborah Hall, and Huan Liu. Modeling temporal patterns of cyberbullying detection with hierarchical attention networks. ACM Transactions on Data Science (TDS), 2(2):8:1–8:23, May 2021. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3441141.

Chakraborty:2020:EDA

[CPAS20] Tanmoy Chakraborty, Noseong Park, Ayush Agarwal, and V. S. Subrahmanian. Ensemble detection and analysis of communities in complex networks. ACM Transactions on Data Science (TDS), 1(1):2:1–2:34, March 2020. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        abs/10.1145/3313374.

Dey:2021:SRC

[DBR21] Paramita Dey, Subhayan Bhattacharya, and Sarbani Roy. A survey on the role of centrality as seed nodes for information propagation in large scale network. ACM Transactions on Data Science (TDS), 2(3):24:1–24:25, August 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

                                        acm.org/doi/10.1145/3465374.

Desmet:2021:RDP

[DC21a] Chance Desmet and Diane J. Cook. Recent developments in privacy-preserving mining of clinical data. ACM Transactions on Data Science (TDS), 2(4):

28:1–28:32, November 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3447774.

Dutta:2021:BDC

[DC21b] Hridoy Sankar Dutta and Tanmoy Chakraborty. Blackmarketdriven collusion on online media: a survey. ACM Transactions on Data Science (TDS), 2(4):43:1–43:??, November 2021. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3517931.

Ding:2021:DPD

[DCZ^{+}21] Xiaofeng Ding, Lin Chen, Pan Zhou, Wenbin Jiang, and Hai Jin. Differentially private deep learning with iterative gradient descent optimization. ACM Transactions on Data Science (TDS), 2(4):34:1–34:27, November 2021. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3491254.

Du:2021:AED

[DLZ^{+}21] Zhekai Du, Jingjing Li, Lei Zhu, Ke Lu, and Heng Tao Shen. Adversarial energy disaggregation. ACM Transactions on Data Science (TDS), 2(4): 27:1–27:16, November 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3477301.

Ding:2020:TAH

[DZWL20] Weilong Ding, Zhuofeng Zhao, Jianwu Wang, and Han Li. Task allocation in hybrid big data analytics for urban IoT applications.

ACM Transactions on Data Science (TDS), 1(3):18:1–18:22, October 2020. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3374751.

Elbassuoni:2020:FSO

[EAYG20] Shady Elbassuoni, Sihem Amer-Yahia, and Ahmad Ghizzawi. Fairness of scoring in online job marketplaces. ACM Transactions on Data Science (TDS), 1(4): 29:1–29:30, December 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3402883.

Fang:2020:HHR

[FZL^{+}20] Liming Fang, Hongwei Zhu, Boqing Lv, Zhe Liu, Weizhi Meng, Yu Yu, Shouling Ji, and Zehong Cao. HandiText: Handwriting recognition based on dynamic characteristics with incremental LSTM. ACM Transactions on Data Science (TDS), 1(4):25:1–25:18, December 2020. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3385189.

Gan:2021:TRM

[GCY^{+}21] Wensheng Gan, Guoting Chen, Hongzhi Yin, Philippe Fournier-Viger, Chien-Ming Chen, and Philip S. Yu. Towards revenue maximization with popular and profitable products. ACM Transactions on Data Science (TDS), 2(4):42:1–42:??, November 2021. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3488058.

Gramaglia:2021:GTP

[GFFS21] Marco Gramaglia, Marco Fiore, Angelo Furno, and Razvan Stanica. GLOVE: Towards privacypreserving publishing of recordlevel-truthful mobile phone trajectories. ACM Transactions on Data Science (TDS), 2(3): 21:1–21:36, August 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3451178.

Gao:2021:MGD

[GLD^{+}21] Hongchao Gao, Yujia Li, Jiao Dai, Xi Wang, Jizhong Han, and Ruixuan Li. Multi-granularity deep local representations for irregular scene text recognition. ACM Transactions on Data Science (TDS), 2(2):15:1–15:18, May 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3446971.

Gan:2020:UMA

[GLZY20] Wensheng Gan, Jerry Chun-Wei Lin, Jiexiong Zhang, and Philip S. Yu. Utility mining across multi-sequences with individualized thresholds. ACM Transactions on Data Science (TDS), 1(2):8:1–8:29, July 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3362070.

Goodman:2021:DBP

[GSM21] Joel Goodman, Sharham Sarkani, and Thomas Mazzuchi. Distancebased probabilistic data augmentation for synthetic minority oversampling. ACM Transactions on Data Science (TDS), 2(4):40:1– 40:??, November 2021. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3510834.

Gao:2021:QTN

[GYZ^{+}21] Yuan Gao, Laurence T. Yang, Dehua Zheng, Jing Yang, and Yaliang Zhao. Quantized tensor neural network. ACM Transactions on Data Science (TDS), 2(4):33:1–33:18, November 2021. CODEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3491255.

Hu:2020:ISI

[HSC20] Haibo Hu, Rik Sarkar, and Zhengzhang Chen. Introduction to the special issue on retrieving and learning from Internet of Things data. ACM Transactions on Data Science (TDS), 1 (4):24:1, December 2020. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3426368.

Iyengar:2021:WDD

[ILI^{+}21] Srinivasan Iyengar, Stephen Lee, David Irwin, Prashant Shenoy, and Benjamin Weil. WattScale: a data-driven approach for energy efficiency analytics of buildings at scale. ACM Transactions on Data Science (TDS), 2 (1):3:1–3:25, January 2021. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3406961.

Jiang:2021:TUH

[JSF^{+}21] Renhe Jiang, Xuan Song, Zipei Fan, Tianqi Xia, Zhaonan Wang, Quanjun Chen, Zekun Cai, and Ryosuke Shibasaki. Transfer urban human mobility via POI embedding over multiple cities. ACM Transactions on Data Science (TDS), 2(1):4:1–4:26, January 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3416914.

Jia:2021:PGM

[JWK^{+}21] Xiaowei Jia, Jared Willard, Anuj Karpatne, Jordan S. Read, Jacob A. Zwart, Michael Steinbach, and Vipin Kumar. Physics-guided machine learning for scientific discovery: an application in simulating lake temperature profiles. ACM Transactions on Data Science (TDS), 2(3):20:1–20:26, May 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3447814.

Li:2020:ISI

[LBZB20] Yanhua Li, Jie Bao, Zhi-Li Zhang, and Saif Benjaafar. Introduction to the special issue on urban computing and smart cities. ACM Transactions on Data Science (TDS), 1(3):14:1–14:2, October 2020. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3412392.

Li:2021:ISI

[LBZB21] Yanhua Li, Jie Bao, Zhi-Li Zhang, and Saif Benjaafar. Introduction to the special issue on urban computing and smart cities.

ACM Transactions on Data Science (TDS), 2(1):2e:1–2e:2, January 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3441679.

Li:2020:GBR

[LCZ^{+}20] Guohui Li, Qi Chen, Bolong Zheng, Hongzhi Yin, Quoc Viet Hung Nguyen, and Xiaofang Zhou. Group-based recurrent neural networks for POI recommendation. ACM Transactions on Data Science (TDS), 1(1):3:1– 3:18, March 2020. CODEN ???? ISSN 2691-1922. URL https:/

/dl.acm.org/doi/abs/10.1145/
3343037.

Li:2020:MAF

[LHJ20] Xiangyang Li, Luis Herranz, and Shuqiang Jiang. Multifaceted analysis of fine-tuning in a deep model for visual recognition. ACM Transactions on Data Science (TDS), 1(1):4:1–4:22, March 2020. CODEN ???? ISSN 2691-

1922. URL https://dl.acm.
org/doi/abs/10.1145/3319500.

Li:2020:PGE

[LKW^{+}20] Yan Li, Pratik Kotwal, Pengyue Wang, Yiqun Xie, Shashi Shekhar, and William Northrop. Physicsguided energy-efficient path selection using on-board diagnostics data. ACM Transactions on Data Science (TDS), 1(3): 22:1–22:28, October 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3406596.

Li:2020:TTR

[LLC^{+}20] Huan Li, Hua Lu, Gang Chen, Ke Chen, Qinkuang Chen, and Lidan Shou. Toward translating raw indoor positioning data into mobility semantics. ACM Transactions on Data Science (TDS), 1(4):26:1–26:37, December 2020. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3385190.

Leng:2021:TEA

[LNP21] Yan Leng, Alejandro Noriega, and Alex Pentland. Tourism event analytics with mobile phone data. ACM Transactions on Data Science (TDS), 2(3):25:1–25:22, August 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3479975.

Liu:2021:SBU

[LNZ21] Xiuming Liu, Edith Ngai, and Dave Zachariah. Scalable belief updating for urban air quality modeling and prediction. ACM Transactions on Data Science (TDS), 2(1):2:1–2:19, January 2021. CODEN ???? ISSN 2691-

1922. URL https://dl.acm.
org/doi/10.1145/3402903.

Liu:2021:DHB

[LWG^{+}21] Yu Liu, Yangtao Wang, Lianli Gao, Chan Guo, Yanzhao Xie, and Zhili Xiao. Deep hash-based relevance-aware data quality assessment for image dark data. ACM Transactions on Data Science (TDS), 2(2):11:1–11:26, May 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

                                        acm.org/doi/10.1145/3420038.

Li:2021:TAN

[LZYZ21] Yamin Li, Jun Zhang, Zhongliang Yang, and Ru Zhang. Topicaware neural linguistic steganography based on knowledge graphs. ACM Transactions on Data Science (TDS), 2(2):10:1–10:13, May 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

                                        acm.org/doi/10.1145/3418598.

Maji:2021:CUF

[MB21] Subhadip Maji and Smarajit Bose. CBIR using features derived by deep learning. ACM Transactions on Data Science (TDS), 2 (3):26:1–26:24, August 2021. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3470568.

Molinaro:2021:SST

[MPS^{+}21] Cristian Molinaro, Chiara Pulice, Anja Subasic, Abigail Bartolome, and V. S. Subrahmanian. STAR: Summarizing Timed Association Rules. ACM Transactions on Data Science (TDS), 2(1): 6:1–6:36, January 2021. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3419107.

Nashaat:2020:AGL

[NGMQ20] Mona Nashaat, Aindrila Ghosh, James Miller, and Shaikh Quader. Asterisk: Generating large training datasets with automatic active supervision. ACM Transactions on Data Science (TDS), 1 (2):13:1–13:25, July 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3385188.

Nashaat:2021:TUR

[NGMQ21] Mona Nashaat, Aindrila Ghosh, James Miller, and Shaikh Quader. TabReformer: Unsupervised representation learning for erroneous data detection. ACM Transactions on Data Science (TDS), 2 (3):18:1–18:29, May 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3447541.

Ooi:2020:IIE

[Ooi20] Beng Chin Ooi. Inaugural issue editorial. ACM Transactions on Data Science (TDS), 1 (1):1:1–1:2, March 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
abs/10.1145/3368254.

Steadman:2021:KDS

[SGJ^{+}21] Liam Steadman, Nathan Griffiths, Stephen Jarvis, Mark Bell, Shaun Helman, and Caroline Wallbank. kD-STR: a method for spatio-temporal data reduction and modelling. ACM Transactions on Data Science (TDS), 2(3):17:1–17:31, May 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3439334.

Song:2021:PPC

[SHC^{+}21] Jie Song, Qiang He, Feifei Chen, Ye Yuan, and Ge Yu. PoBery: Possibly-complete big data queries with probabilistic data placement and scanning. ACM Transactions on Data Science (TDS), 2(3):23:1–23:28, August 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3465375.

Sainju:2020:MRS

[SJ20] Arpan Man Sainju and Zhe Jiang. Mapping road safety features from streetview imagery: a deep learning approach. ACM Transactions on Data Science (TDS), 1(3): 15:1–15:20, October 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3362069.

Shi:2021:NAT

[SKRR21] Tian Shi, Yaser Keneshloo, Naren Ramakrishnan, and Chandan K. Reddy. Neural abstractive text summarization with sequence-to-sequence models. ACM Transactions on Data Science (TDS), 2 (1):1:1–1:37, January 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3419106.

Schmeisser:2020:CCS

[SS20] Stephan Schmeißer and Gregor Schiele. coSense: The collaborative sensing middleware for the Internet-of-things. ACM Transactions on Data Science (TDS), 1(4):28:1–28:21, December 2020. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3395233.

Tian:2020:UEB

[TLL^{+}20] Zhihong Tian, Chaochao Luo, Hui Lu, Shen Su, Yanbin Sun, and Man Zhang. User and entity behavior analysis under urban big data. ACM Transactions on Data Science (TDS), 1(3): 16:1–16:19, October 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3374749.

Theil:2020:EFU

[TSS20] Christoph Kilian Theil, Sanja Stajner, and Heiner Stuckenschmidt. Explaining financial uncertainty through specialized word embeddings. ACM Transactions on Data Science (TDS), 1(1):6:1–6:19, March 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
abs/10.1145/3343039.

Tian:2021:TST

[TZW^{+}21] Xiancai Tian, Baihua Zheng, Yazhe Wang, Hsiao-Ting Huang, and Chih-Chieh Hung. TRIPDE-CODER: Study travel time attributes and route preferences of metro systems from smart card data. ACM Transactions on Data Science (TDS), 2(3):16:1–16:21, August 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3430768.

Wu:2020:DST

[WC20] Gene P. K. Wu and Keith C. C.

Chan. Discovery of spatiotemporal patterns in multivariate spatial time series. ACM Transactions on Data Science (TDS), 1(2):11:1–11:22, July 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3374748.

Wang:2021:TBD

[WGL^{+}21] Chaoyang Wang, Zhiqiang Guo, Jianjun Li, Guohui Li, and Peng Pan. A text-based deep reinforcement learning framework using self-supervised graph representation for interactive recommendation. ACM Transactions on Data Science (TDS), 2(4): 44:1–44:??, November 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3522596.

Ward:2020:PAA

[WLM20] Katrina Ward, Dan Lin, and Sanjay Madria. A parallel algorithm for anonymizing large-scale trajectory data. ACM Transactions on Data Science (TDS), 1 (1):5:1–5:26, March 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
abs/10.1145/3368639.

Xu:2020:AWF

[XFW^{+}20] Xin Xu, Yanjie Fu, Jingyi Wu, Yuqi Wang, Zeyu Huang, Zhiguo Fu, and Minghao Yin. Adaptive weighted finite mixture model: Identifying the feature-influence of real estate. ACM Transactions on Data Science (TDS), 1(3): 20:1–20:16, October 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3379560.

Xia:2021:SBA

[XGB^{+}21] Feng Xia, Teng Guo, Xiaomei Bai, Adrian Shatte, Zitao Liu, and Jiliang Tang. SUMMER: Bias-aware prediction of graduate employment based on educational big data. ACM Transactions on Data Science (TDS), 2(4):39:1–39:??, November 2021. CODEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3510361.

Xie:2020:UFR

[XS20] Yiqun Xie and Shashi Shekhar.

A unified framework for robust and efficient hotspot detection in smart cities. ACM Transactions on Data Science (TDS), 1(3): 17:1–17:29, October 2020. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3379562.

Yang:2021:BRP

[YDZH21] Ru Yang, Yuhui Deng, Yi Zhou, and Ping Huang. Boosting the restoring performance of deduplication data by classifying backup metadata. ACM Transactions on Data Science (TDS), 2(2): 13:1–13:16, May 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3437261.

Yu:2020:MBI

[YLY20] Lei Yu, Guohui Li, and Ling Yuan. Maximizing boosted influence spread with edge addition in online social networks. ACM Transactions on Data Science (TDS), 1(2):9:1–9:21, July 2020. CODEN ???? ISSN 2691-1922. URL https://dl.

                                        acm.org/doi/10.1145/3364993.

Yang:2020:MFR

[YXL20] Luoying Yang, Zhou Xu, and Jiebo Luo. Measuring female representation and impact in films over time. ACM Transactions on Data Science (TDS), 1(4): 30:1–30:14, December 2020. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3411213.

Yang:2020:SAT

[YZL^{+}20] Zhong Yang, Bolong Zheng, Guohui Li, Nguyen Quoc Viet Hung, Guanfeng Liu, and Kai Zheng. Searching activity trajectories by exemplar. ACM Transactions on Data Science (TDS), 1(3): 19:1–19:18, October 2020. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3379561.

Zandavi:2021:MUR

[ZCA21] Seid Miad Zandavi, Vera Chung, and Ali Anaissi. Multi-user remote lab: Timetable scheduling using simplex nondominated sorting genetic algorithm. ACM Transactions on Data Science (TDS), 2(2):14:1–14:13, May 2021. CODEN ???? ISSN 2691-

                                        1922. URL https://dl.acm.
                                        org/doi/10.1145/3437260.

Zoppi:2021:UAD

[ZCCB21] Tommaso Zoppi, Andrea Ceccarelli, Tommaso Capecchi, and Andrea Bondavalli. Unsupervised anomaly detectors to detect intrusions in the current threat landscape. ACM Transactions on Data Science (TDS), 2(2):7:1–7:26, May 2021. CO-DEN ???? ISSN 2691-1922.

URL https://dl.acm.org/doi/
10.1145/3441140.

Zhang:2021:SIR

[ZLJC21] Xinyu Zhang, Xiaocui Li, Xiao-Yuan Jing, and Li Cheng. Simultaneous image reconstruction and feature learning with 3D-CNNs for image set-based classification. ACM Transactions on Data Science (TDS), 2(2):12:1–12:13, May 2021. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3420037.

Zeidan:2020:GEL

[ZLZ^{+}20] Ayman Zeidan, Eemil Lagerspetz, Kai Zhao, Petteri Nurmi, Sasu Tarkoma, and Huy T. Vo. Geo-Match: Efficient large-scale map matching on Apache Spark. ACM Transactions on Data Science (TDS), 1(3):21:1–21:30, October 2020. CODEN ???? ISSN 2691-1922. URL https://dl.

acm.org/doi/10.1145/3402904.

Zhang:2021:HSS

[ZMS^{+}21] Jiaru Zhang, Ruhui Ma, Tao Song, Yang Hua, Zhengui Xue, Chenyang Guan, and Haibing Guan. Hierarchical satellite system graph for approximate nearest neighbor search on big data. ACM Transactions on Data Science (TDS), 2(4):32:1– 32:15, November 2021. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3488377.

Zhou:2021:ISI

[ZS21] Ke Zhou and Jingkuan Song. Introduction to the special issue on learning-based support for data science applications. ACM Transactions on Data Science (TDS), 2(2):9:1, May 2021. CO-DEN ???? ISSN 2691-1922.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3450751.